-2x - 15 = -9 (x + 4)
Answer by stanbon(48551)

-2x - 15 = -9 (x + 4)
-2x - 15 = -9x - 36
Add 9x to both sides: Add 15 to both sides
7x = -21
x = -3
============
Cheers,
Stan H.